OpenStudy (anonymous):

Name the property the equation illustrates. a + c + 2b = a + 2b + c Identity Property of Addition Distributive Property Associative Property of Addition Commutative Property of Addition

OpenStudy (anonymous):

Commutative Property http://www.purplemath.com/modules/numbprop.htm

OpenStudy (anonymous):

can you help me with some more?

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 21 vowels and 15 consonants, what is the probability you will choose a consonant and then a vowel? 35/8 1/36 35/4 35/144

OpenStudy (anonymous):

\[\frac{15}{36}*\frac{21}{36}=\frac{35}{144} \]